how do you find the equation of the quadratic function. it has a graph with x-intercepts 2 and -1 and y-intercepts 6
Answer by funmath(2933)

how do you find the equation of the quadratic function. it has a graph with x-intercepts 2 and -1 and y-intercepts 6
y=a(x-m)(x-n) where m and n are the x-intercepts.
y=a(x-2)(x-(-1)) Plug in the x-intercepts.
y=a(x-2)(x+1) The y-intercept is (0,6), let x=0 and y=6 and solve for a
6=a(0-2)(0+1)
6=a(-2)(1)
6=-2a
6/-2=-2a/-2
-3=a
Here's the equation that you're looking for:
y=-3(x-2)(x+1) <--most teachers are happy here, but if you have to multiply it out:
y=-3(x^2+x-2x-2)
y=-3(x^2-x-2)
y=-3x^2+3x+6
:
Happy Calculating!!!